OpenStudy (turingtest):

Easier challenge: Derivatives - Is the power rule simple? 1)Prove the power rule for integral exponents. 2)Given that result, prove that the power rule holds for all rational exponents. 3)Using any known derivative formulas/techniques, prove the power rule holds for irrational numbers as well.

OpenStudy (kinggeorge):

This could take a while to write out...

OpenStudy (turingtest):

the first one, yeah... I'll accept shorthand :)

OpenStudy (turingtest):

you may accept the chain rule as given...

OpenStudy (kinggeorge):

Basically, if we can show it for \(x^n\), by the properties of derivatives, we'll have proved it for any polynomial \[a_0+a_1x+a_2x^2+...+a_nx^n\]So let's try and show it for \(x^n\)!

OpenStudy (turingtest):

really? new one on me...

OpenStudy (experimentx):

power rule can be proved via inductive method using product rule for positive integers.

OpenStudy (turingtest):

the first one must be proved from first principles after that accept the chain rule, and ... no more hints besides, I wanna see different ways

OpenStudy (kinggeorge):

\[\lim_{h \rightarrow 0} {f(x+h)-f(x) \over h}=\lim_{h \rightarrow 0} {(x+h)^n-x^n \over h}\]If we expand the numerator, we get\[=\binom{n}{0}x^n+\binom{n}{1}x^{n-1}h+\binom{n}{2}x^{n-2}h^2+...+\binom{n}{n}h^n-x^n\]From this, we can cancel the \(x^n\) terms.Then, we can cancel an h from every term This leaves us with\[\lim_{h \rightarrow 0} {nx^{n-1} +h(\text{a bunch of stuff})}\]Which is equal to zero since all the "stuff" goes to 0. This leaves us with the derivative \(nx^{n-1}\). That proves 1 from the definition of derivative without any extra assumptions.

OpenStudy (kinggeorge):

This proves it for any polynomial \(a_0+a_1x+...+a_nx^n\) since we can split it apart into \(n+1\) separate polynomials and take the derivative of each, and then add them back together. We can also factor constants out of limits, so we're good there.

OpenStudy (turingtest):

yes :) now just assuming the chain rule and do part 2)

OpenStudy (kinggeorge):

For rational exponents, we need to prove it for \[x^{1/a}\]first. Then we can simply raise this to the \(n\)-th power.

OpenStudy (turingtest):

not quite the way I know it, so feel free to demonstrate

OpenStudy (kinggeorge):

I've never done this before, so I'm winging it here.

OpenStudy (turingtest):

you may request a hint if you like

OpenStudy (anonymous):

\[x^r=e^{r\ln(x)}\]

OpenStudy (turingtest):

you are doing problem 3) satellite you can't assume to know the derivative of e^r until part 3) at least that's the way I want to do it, but your way would work of course

OpenStudy (anonymous):

not to be picky but if say \(\pi\) is irrational, how do we define \(x^{\pi}\) for real values of x?

OpenStudy (turingtest):

I dunno

OpenStudy (anonymous):

here is one i like prove \((fg)'=f'g+g'f\) using only that \((f^2)'=2ff'\)

OpenStudy (kinggeorge):

I'm gonna request a small hint for #2 here.

OpenStudy (anonymous):

#2 use chain rule

OpenStudy (turingtest):

\[y=x^{n/m}\implies y^m=x^n\]

OpenStudy (kinggeorge):

Is implicit differentiation allowed?

OpenStudy (turingtest):

yes, that's what I meant by chain rule

OpenStudy (kinggeorge):

Well then, \[y^m=x^n \Longrightarrow my^{m-1}\;\;dy=nx^{n-1}\;\;dx\]So \[{dy \over dx}={nx^{n-1} \over my^{m-1}}\]Then, since \(y=x^{n/m}\), we know that \[\Large y^{m-1}=(n^{n/m})^{m-1}=x^{{nm \over m}-{n \over m}}=x^{n-{n \over m}}\]If we plug that into our equation for \({dy/dx}\), we have \[\Large {dy \over dx}={nx^{n-1} \over mx^{n-{n \over m}}}\]Which in turn equals \[\Large {n \over m}\cdot x^{(n-1)-(n-{n \over m})} ={n \over m} \cdot x^{{n \over m} -1}\]Which is what we wanted to prove.

OpenStudy (turingtest):

haha, indeed! nicely done :) (wish I could give 2 medals) the algebra is the only pain in the butt once you know what to do

OpenStudy (turingtest):

part 3) is actually the fastest to prove you may use all common derivative knowledge at your disposal for that one (chain rule, well-known derivatives and techniques, etc.)

OpenStudy (kinggeorge):

Openstudy just crashed for me. Give me a minute to write it all up again.

OpenStudy (turingtest):

I hate when that happens :( no rush...

OpenStudy (turingtest):

I copy what I have regularly to avoid such catastrophes

OpenStudy (kinggeorge):

Using satellite's suggestion, \[\large y=x^a =e^{a \ln(x)} \Longrightarrow {dy \over dx}=e^{a \ln(x)} \cdot {a \over x}\]Simplifying this, we have \[{dy \over dx} ={ax^a \over x}=ax^{a-1}\]

OpenStudy (kinggeorge):

That was an excellent exercise in the fundamentals of calculus.

OpenStudy (turingtest):

tadah! 3/3, nice job :) I'm glad you enjoyed it (this is directly from the MIT calc I lectures, btw)

OpenStudy (turingtest):

I will now post a bonus problem, which I have posted before, but did not receive a satisfactory explanation to. Let those who don't know the trick attempt it.